# When LLMs Play the Telephone Game: Cumulative Changes and Attractors in Iterated Cultural Transmissions

### v2

@@ Appendix A Additional details on the methods
Data analyses revealed that, on the Continue task, when using Mistral-7B, agents of the chains would sometimes start outputting very long text by filling them with “#some_keyword”. As this behavior created a few outliers, we thought it would be better to filter-out those “#some_keyword” when performing the main analyses. This behavior is nevertheless an interesting result, reminiscent of the collapsing dynamics found when training LLMs on their own output [shumailov2023curse]. We therefore discuss it separately in Appendix B.
@@
  • Positivity. We employ the SentimentIntensityAnalyzer tool from the NLTK library to assess the positivity of generated texts. The tool is based on the Valence Aware Dictionary and sEntiment Reasoner (VADER) method [hutto2014vader], which is a lexicon and rule-based sentiment analysis tool specifically designed for social media data. It uses a combination of lexical features, such as words and their semantic orientation, to determine the overall sentiment of a text. In the VADER method, every word in the vocabulary is rated with respect to its positive or negative sentiment and the intensity of that sentiment. The SentimentIntensityAnalyzer uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
@@ Appendix B Additional figures and analyses
Statistical models
@@
We performed statistical analyses using the Python package pymc [wiecki_pymc-devspymc_2024] to fit Bayesian models.
@@
To determine the significance of the difference between estimated parameters, we computed the 95% credibility intervals of the difference by sampling from the posteriors. In Figures S3 and S4, we provide the matrices indicating the mean of this difference between estimated parameters. Stars indicate that the credibility interval does not contain 0.
Figure S3: Differences between posterior estimates for the statistical model predicting attractor Strength. Differences between estimate for the effect of Model (left), Measure (center) and Task (right) on attractor Strength. Values and color represent the mean value of the 95% credibility intervals. Stars indicate that this interval does not contain 0.
Figure S4: Differences between posterior estimates for the statistical model predicting attractor Position. Differences between estimate for the effect of Model (top row) and Task (bottom row) on attractor Position for toxicity (first column), positivity (second column), difficulty (third column) and length (fourth column). Values and color represent the mean value of the 95% credibility intervals. Stars indicate that this interval does not contain 0.
Discontinuities and collapsing behavior
@@
Figure S5 shows the average length of text generated with the Mistral-7B model chain on the Continue task for five different seeds of the same story. We can observe several discontinuities in terms of the generated text length, i.e. at some iterations the length drastically increases or decreases. It is interesting to note that when the length decreases, it returns to the original value as before the first discontinuity. This suggests the existence of an attractor regarding this specific length. To better understand the cause of these discontinuities, Figure 2 shows examples of stories generated before and after those discontinuities (for the seed number three in figure S5). We can see that at generation 14 the model abruptly starts to generate many hashtags. It generates 283 hashtags, compared to 12 in the previous generation. At generation 45, we can see that the overall quality of the text decreased into generating solely hashtags and brief descriptions. This reduction in text quality is reminiscent of collapsing dynamics observed in iterative chains of LLMs, where each model was trained on the output of a previous one [shumailov2023curse].

### v3

@@ Appendix A Additional details on the methods
Data analyses revealed that, on the Continue task, when using Mistral-7B, agents of the chains would sometimes start outputting very long text by filling them with “#some_keyword”. As this behavior created a few outliers, we thought it would be better to filter-out those “#some_keyword” when performing the main analyses. This behavior is nevertheless an interesting result, reminiscent of the collapsing dynamics found when training LLMs on their own output [Shumailov et al., 2023]. We therefore discuss it separately in Appendix B.
@@
  • Positivity. We employ the SentimentIntensityAnalyzer tool from the NLTK library to assess the positivity of generated texts. The tool is based on the Valence Aware Dictionary and sEntiment Reasoner (VADER) method [Hutto and Gilbert, 2014], which is a lexicon and rule-based sentiment analysis tool specifically designed for social media data. It uses a combination of lexical features, such as words and their semantic orientation, to determine the overall sentiment of a text. In the VADER method, every word in the vocabulary is rated with respect to its positive or negative sentiment and the intensity of that sentiment. The SentimentIntensityAnalyzer uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
@@ B.1 Fitted Linear regressions
B.2 Robustness check and controls
B.2.1 Increasing the size of the set of initial texts
To verify that the results presented in the main text hold when using a larger set of initial texts, we ran additional experiments with Meta-Llama-3-8b using 100 different initial texts, using the same sampling procedure as for the main experiment (see Section A). We observe that the trends remain the same with 100 initial texts as with 20 initial texts, and that the standard deviation intervals for 100 initial texts include – or are very close to including – the values estimated with 20 initial texts (with only one exception out of 24 values (Figure S3)
Figure S3: Attractor strenghts and positions over 100 initial texts Attractor position (first row) and strength (second row) for four text properties (column) and three tasks (colors). The height of the bars are the position and strength of the attractors for a set of 100 initial texts. The red crosses indicate the values estimated in the main experiment using a set of 20 initial texts (Section 4.3). This reveals that our results are robust to increasing the sample size of the initial texts dataset.
B.2.2 Different phrasing of the instructions prompts
To verify that the results presented in the main text hold when using different phrasings of the initial prompts, we ran additional experiments with Llama-3-8B using 5 different phrasings of instructions prompts for each of the three. Those exact prompts are as follows:
Rephrase
  • "You will be given a text. Your job is to reword this text without changing its meaning. Only provide your revised text, nothing else. Here is the text:"
  • "A text will be provided to you. Your task is to rephrase it, keeping its meaning intact. Only output the rephrased text, and nothing additional. Here is the text:"
  • "You’ll receive a text, and your job is to rephrase it without altering its meaning. Just output your new version, nothing more. Here is the text:"
  • "A text will be sent to you. Your role is to rephrase it while keeping the meaning the same. Only display your new text, with nothing extra. Here is the text:"
  • "You will be provided with a text. Your task is to reword it without changing the intended meaning. Output just your rephrased text, nothing else. Here is the text:"
Inspiration
  • "You’ll be given a text, and your task is to craft a new, original text inspired by it. Only provide your new version, with nothing additional. Here is the text:"
  • "A text will be provided to you. Your task is to create an original text based on this inspiration. Just output your new version, nothing extra. Here is the text:"
  • "You will receive a text. Your job is to generate an original text inspired by it. Only show your new version, without adding anything else. Here is the text:"
  • "You’ll receive a text, and your task is to create a new text inspired by it. Simply display your new version, with no additional output. Here is the text:"
  • "You will be sent a text. Your role is to produce an original text inspired by it. Only present your new text, nothing more. Here is the text:"
Continue
  • "You’ll be provided with a text. Your job is to extend this text. Only output your continuation, nothing additional. Here is the text:"
  • "A text will be given to you. Your task is to carry on from this text. Just display your new continuation, with nothing extra. Here is the text:"
  • "You’ll receive a text. Your role is to continue from where this text ends. Only present your extension, with no additional output. Here is the text:"
  • "A text will be sent to you, and your job is to complete it by continuing from its end. Only output your continuation, nothing more. Here is the text:"
  • "You will get a text. Your task is to extend it further. Just display your continuation, without adding anything else. Here is the text:"
This revealed that our results are robust to different paraphrasing of the same prompt: indeed, the interval of the standard deviations over the 5 paraphrased prompts always contains – or almost contains – the values estimated with the original prompt in the main experiment. (Figure S4)
Figure S4: Attractor strenghts and positions for different phrasings of the intruction prompts Attractor position (first row) and strength (second row) for four text properties (column) and three tasks (colors). The height of the bars represent the average position and strength of the attractors over the five different phrasings. Individual dots correspond to each different phrasing. The red crosses indicate the values estimated in the main experiment using a set of 20 initial texts (Section 4.3). This reveals that our results are robust to increasing the sample size of the initial texts dataset.
B.2.3 Effect of less versus more constrained tasks on attractors
In the results from the main experiments, we observed that less constrained tasks (such as Continue) lead to stronger attractors than more constrained tasks (such as Rephrase). However, these two tasks vary on other dimensions than only the room for variation allowed. To verify our hypothesis, we conducted additional experiments with Llama-3-8B where we instructed the LLM to create a new text by making either “very minor”, “a few small”, “moderate”, “notable” or “radical” changes to the received text.
The exact prompts were as follows:
  • "You will receive a text. Your task is to create a new original text by taking inspiration from this text, making only very minor changes. Just output your new text, nothing else. Here is the text:"
  • "You will receive a text. Your task is to create a new original text by taking inspiration from this text, making a few small changes. Just output your new text, nothing else. Here is the text:"
  • "You will receive a text. Your task is to create a new original text by taking inspiration from this text, making moderate changes. Just output your new text, nothing else. Here is the text:"
  • "You will receive a text. Your task is to create a new original text by taking inspiration from this text, making notable changes. Just output your new text, nothing else. Here is the text:"
  • "You will receive a text. Your task is to create a new original text by taking inspiration from this text, making radical changes. Just output your new text, nothing else. Here is the text:"
This experiment confirmed our initial interpretation, as less constrained tasks indeed exhibit stronger attractors (FigureS5). The only exception is Length. We interpret this by the fact that going from “very minor” to “radical” changes mainly refers to the semantic of the text. As a consequence, the instructions to make “minor changes” is not more constrained than to make “radical changes” with respect to Length.
Figure S5: Effect of less versus more constrained tasks on attractors Attractor position (first row) and strength (second row) for four text properties (column) as a function of the room for variation allowed in the instruction. These results confirm our hypothesis that less constrained tasks lead to stronger attractors. The only exception is Length. We interpret this by the fact that going from “very minor” to “radical” changes mainly refers to the semantic of the text. As a consequence, the instructions to make “minor changes” is not more constrained than to make “radical changes” with respect to Length.
B.2.4 Effect of heterogeneous transmission chains
As the study of LLMs’ multi-turn dynamics remains unexplored, it appeared necessary to start with a setting where causal variables can be isolated, so as to lay the foundation for incrementally improving our understanding of LLMs cultural dynamics. As a consequence, the main experiments focused on homogeneous chains where the multi-turn dynamics of different models can be assessed in isolation. However, we also investigate the effect of having heterogeneous chains, where two different models interact (Figure S6). We compared 4 conditions: homogeneous chains of Mistral-7B models, homogeneous chains of Llama-3-8B models, heterogeneous chains of Mistral-7B and Llama-3-8B models starting with Mistral-7B and heterogeneous chains of Mistral-7B and Llama-8B models starting with Llama-3-8B. We found that although the position of the attractor in heterogeneous chains is often in between the position of the homogeneous chains’ attractors, it is not systematically the case.
Figure S6: Comparison of homogeneous and heterogeneous transmission chains Attractor position (first row) and strength (second row) for four text properties (column) and three different tasks, for homogeneous chains of Llama-3-8B models (blue), homogeneous chains of Mistral-7B models (red), heterogeneous chains of Mistral-7B and Llama-3-8B models starting with Mistral-7B (dark purple) and heterogeneous chains of Mistral-7B and Llama-8B models starting with Llama-3-8B (light purple).
B.3 Statistical models
B.3.1 Kolmogorov-Smirnov tests
To quantitatively evaluate whether multi-turn transmission lead to significantly different outcomes than single-turn transmission, we use Kolmogorov–Smirnov (KS) tests [Massey, 1951] to estimate the compare distributions of text properties after a single interaction and after multiple interactions. Although we calculate p-values repeatedly (once for each generation, model and task), in our cases applying corrections to control for the False Discovery Rate (FDR) was in our case not necessary. Indeed, controlling for the FDR using methods such as Bonferroni correction is necessary in cases where many statistical tests are conducted, as it increases the likelihood that at least one of them is found significant “by chance” (Family-Wise Error Rate; Colas, 2022). With respect to our set-up, this would have been necessary if we concluded that multi-turn interactions lead to significant differences when at least one of the 49 p-values (one by generation) was below the significance threshold. However, we rather report all p-values, and draw conclusions from the trend observed over all generations, rather than from single p-values. Therefore, if we observe a general trend of decreasing p-values, we can conclude that observing the given distribution under the null hypothesis becomes less and less likely with generations. FDR corrections are thus unnecessary for drawing this conclusion.
B.3.2 Bayesian models
@@
We performed statistical analyses using the Python package pymc [Wiecki et al., 2024] to fit Bayesian models.
@@
To determine the significance of the difference between estimated parameters, we computed the 95% credibility intervals of the difference by sampling from the posteriors. In Tables 4 to 12, we provide those credibility intervals. Intervals that do not contain 0 signal statistically significant differences.
rephrase inspiration continue
rephrase [0.0000 ; 0.0000] [-0.2372 ; -0.0353] [-0.4187 ; -0.2195]
inspiration [0.0353 ; 0.2372] [0.0000 ; 0.0000] [-0.2824 ; -0.0840]
continue [0.2195 ; 0.4187] [0.0840 ; 0.2824] [0.0000 ; 0.0000]
Table 2: 95% Credible Intervals for posterior differences between prompts for attractor strength
gpt-4o-mini gpt-3.5-turbo-0125 Meta-Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2 Meta-Llama-3-70B-Instruct Mixtral-8x7B-Instruct-v0.1
gpt-4o-mini [0.0000 ; 0.0000] [-0.3539 ; -0.0699] [-0.3289 ; -0.0451] [-0.2361 ; 0.0487] [-0.1644 ; 0.1208] [-0.3501 ; -0.0649]
gpt-3.5-turbo-0125 [0.0699 ; 0.3539] [0.0000 ; 0.0000] [-0.1161 ; 0.1673] [-0.0228 ; 0.2603] [0.0490 ; 0.3346] [-0.1385 ; 0.1478]
Meta-Llama-3-8B-Instruct [0.0451 ; 0.3289] [-0.1673 ; 0.1161] [0.0000 ; 0.0000] [-0.0500 ; 0.2351] [0.0256 ; 0.3087] [-0.1617 ; 0.1209]
Mistral-7B-Instruct-v0.2 [-0.0487 ; 0.2361] [-0.2603 ; 0.0228] [-0.2351 ; 0.0500] [0.0000 ; 0.0000] [-0.0686 ; 0.2134] [-0.2561 ; 0.0277]
Meta-Llama-3-70B-Instruct [-0.1208 ; 0.1644] [-0.3346 ; -0.0490] [-0.3087 ; -0.0256] [-0.2134 ; 0.0686] [0.0000 ; 0.0000] [-0.3282 ; -0.0430]
Mixtral-8x7B-Instruct-v0.1 [0.0649 ; 0.3501] [-0.1478 ; 0.1385] [-0.1209 ; 0.1617] [-0.0277 ; 0.2561] [0.0430 ; 0.3282] [0.0000 ; 0.0000]
Table 3: 95% Credible Intervals for posterior differences between model for attractor strength
toxicity positivity difficulty length
toxicity [0.0000 ; 0.0000] [0.1345 ; 0.3626] [0.1274 ; 0.3584] [0.2422 ; 0.4747]
positivity [-0.3626 ; -0.1345] [0.0000 ; 0.0000] [-0.1201 ; 0.1094] [-0.0054 ; 0.2273]
difficulty [-0.3584 ; -0.1274] [-0.1094 ; 0.1201] [0.0000 ; 0.0000] [-0.0014 ; 0.2330]
length [-0.4747 ; -0.2422] [-0.2273 ; 0.0054] [-0.2330 ; 0.0014] [0.0000 ; 0.0000]
Table 4: 95% Credible Intervals for posterior differences between measures for attractor strength
gpt-4o-mini gpt-3.5-turbo-0125 Meta-Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2 Meta-Llama-3-70B-Instruct Mixtral-8x7B-Instruct-v0.1
gpt-4o-mini [0.0000 ; 0.0000] [-0.0012 ; 0.0069] [-0.0033 ; 0.0049] [-0.0016 ; 0.0064] [-0.0043 ; 0.0037] [-0.0010 ; 0.0070]
gpt-3.5-turbo-0125 [-0.0069 ; 0.0012] [0.0000 ; 0.0000] [-0.0061 ; 0.0020] [-0.0045 ; 0.0035] [-0.0071 ; 0.0009] [-0.0039 ; 0.0042]
Meta-Llama-3-8B-Instruct [-0.0049 ; 0.0033] [-0.0020 ; 0.0061] [0.0000 ; 0.0000] [-0.0024 ; 0.0056] [-0.0051 ; 0.0030] [-0.0019 ; 0.0063]
Mistral-7B-Instruct-v0.2 [-0.0064 ; 0.0016] [-0.0035 ; 0.0045] [-0.0056 ; 0.0024] [0.0000 ; 0.0000] [-0.0066 ; 0.0014] [-0.0034 ; 0.0046]
Meta-Llama-3-70B-Instruct [-0.0037 ; 0.0043] [-0.0009 ; 0.0071] [-0.0030 ; 0.0051] [-0.0014 ; 0.0066] [0.0000 ; 0.0000] [-0.0007 ; 0.0073]
Mixtral-8x7B-Instruct-v0.1 [-0.0070 ; 0.0010] [-0.0042 ; 0.0039] [-0.0063 ; 0.0019] [-0.0046 ; 0.0034] [-0.0073 ; 0.0007] [0.0000 ; 0.0000]
Table 5: 95% Credible Intervals for posterior differences between model for attractor position - toxicity
rephrase inspiration continue
rephrase [0.0000 ; 0.0000] [-0.0031 ; 0.0000] [-0.0040 ; -0.0009]
inspiration [-0.0000 ; 0.0031] [0.0000 ; 0.0000] [-0.0025 ; 0.0006]
continue [0.0009 ; 0.0040] [-0.0006 ; 0.0025] [0.0000 ; 0.0000]
Table 6: 95% Credible Intervals for posterior differences between prompt for attractor position - toxicity
gpt-4o-mini gpt-3.5-turbo-0125 Meta-Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2 Meta-Llama-3-70B-Instruct Mixtral-8x7B-Instruct-v0.1
gpt-4o-mini [0.0000 ; 0.0000] [-0.0982 ; 0.3959] [0.2023 ; 0.6959] [-0.1887 ; 0.3048] [-0.0172 ; 0.4735] [-0.1299 ; 0.3673]
gpt-3.5-turbo-0125 [-0.3959 ; 0.0982] [0.0000 ; 0.0000] [0.0521 ; 0.5498] [-0.3395 ; 0.1533] [-0.1712 ; 0.3262] [-0.2771 ; 0.2166]
Meta-Llama-3-8B-Instruct [-0.6959 ; -0.2023] [-0.5498 ; -0.0521] [0.0000 ; 0.0000] [-0.6410 ; -0.1471] [-0.4723 ; 0.0229] [-0.5813 ; -0.0818]
Mistral-7B-Instruct-v0.2 [-0.3048 ; 0.1887] [-0.1533 ; 0.3395] [0.1471 ; 0.6410] [0.0000 ; 0.0000] [-0.0767 ; 0.4176] [-0.1860 ; 0.3139]
Meta-Llama-3-70B-Instruct [-0.4735 ; 0.0172] [-0.3262 ; 0.1712] [-0.0229 ; 0.4723] [-0.4176 ; 0.0767] [0.0000 ; 0.0000] [-0.3566 ; 0.1401]
Mixtral-8x7B-Instruct-v0.1 [-0.3673 ; 0.1299] [-0.2166 ; 0.2771] [0.0818 ; 0.5813] [-0.3139 ; 0.1860] [-0.1401 ; 0.3566] [0.0000 ; 0.0000]
Table 7: 95% Credible Intervals for posterior differences between model for attractor position - positivity
rephrase inspiration continue
rephrase [0.0000 ; 0.0000] [-0.4662 ; -0.1133] [-0.4465 ; -0.0938]
inspiration [0.1133 ; 0.4662] [0.0000 ; 0.0000] [-0.1570 ; 0.1913]
continue [0.0938 ; 0.4465] [-0.1913 ; 0.1570] [0.0000 ; 0.0000]
Table 8: 95% Credible Intervals for posterior differences between prompt for attractor position - positivity
gpt-4o-mini gpt-3.5-turbo-0125 Meta-Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2 Meta-Llama-3-70B-Instruct Mixtral-8x7B-Instruct-v0.1
gpt-4o-mini [0.0000 ; 0.0000] [-2.6049 ; 2.8520] [-2.9991 ; 2.4390] [-2.6157 ; 2.7268] [-2.9831 ; 2.4137] [-2.6832 ; 2.7637]
gpt-3.5-turbo-0125 [-2.8520 ; 2.6049] [0.0000 ; 0.0000] [-3.1499 ; 2.2758] [-2.7682 ; 2.6207] [-3.1053 ; 2.2984] [-2.8099 ; 2.6017]
Meta-Llama-3-8B-Instruct [-2.4390 ; 2.9991] [-2.2758 ; 3.1499] [0.0000 ; 0.0000] [-2.3698 ; 3.0545] [-2.7175 ; 2.7392] [-2.4107 ; 3.0411]
Mistral-7B-Instruct-v0.2 [-2.7268 ; 2.6157] [-2.6207 ; 2.7682] [-3.0545 ; 2.3698] [0.0000 ; 0.0000] [-3.0124 ; 2.3806] [-2.6899 ; 2.6762]
Meta-Llama-3-70B-Instruct [-2.4137 ; 2.9831] [-2.2984 ; 3.1053] [-2.7392 ; 2.7175] [-2.3806 ; 3.0124] [0.0000 ; 0.0000] [-2.3859 ; 3.0178]
Mixtral-8x7B-Instruct-v0.1 [-2.7637 ; 2.6832] [-2.6017 ; 2.8099] [-3.0411 ; 2.4107] [-2.6762 ; 2.6899] [-3.0178 ; 2.3859] [0.0000 ; 0.0000]
Table 9: 95% Credible Intervals for posterior differences between model for attractor position - difficulty
rephrase inspiration continue
rephrase [0.0000 ; 0.0000] [-3.0072 ; 2.2777] [-2.6556 ; 2.6528]
inspiration [-2.2777 ; 3.0072] [0.0000 ; 0.0000] [-2.2601 ; 2.9840]
continue [-2.6528 ; 2.6556] [-2.9840 ; 2.2601] [0.0000 ; 0.0000]
Table 10: 95% Credible Intervals for posterior differences between prompt for attractor position - difficulty
gpt-4o-mini gpt-3.5-turbo-0125 Meta-Llama-3-8B-Instruct Mistral-7B-Instruct-v0.2 Meta-Llama-3-70B-Instruct Mixtral-8x7B-Instruct-v0.1
gpt-4o-mini [0.0000 ; 0.0000] [-1.8746 ; 3.6795] [-1.8366 ; 3.7029] [-2.1079 ; 3.4919] [-2.2197 ; 3.3266] [-2.0495 ; 3.4024]
gpt-3.5-turbo-0125 [-3.6795 ; 1.8746] [0.0000 ; 0.0000] [-2.7134 ; 2.8132] [-2.9561 ; 2.5396] [-3.1474 ; 2.4007] [-2.9893 ; 2.5194]
Meta-Llama-3-8B-Instruct [-3.7029 ; 1.8366] [-2.8132 ; 2.7134] [0.0000 ; 0.0000] [-3.0299 ; 2.5205] [-3.1786 ; 2.3807] [-3.0437 ; 2.4675]
Mistral-7B-Instruct-v0.2 [-3.4919 ; 2.1079] [-2.5396 ; 2.9561] [-2.5205 ; 3.0299] [0.0000 ; 0.0000] [-2.9116 ; 2.6355] [-2.7884 ; 2.7556]
Meta-Llama-3-70B-Instruct [-3.3266 ; 2.2197] [-2.4007 ; 3.1474] [-2.3807 ; 3.1786] [-2.6355 ; 2.9116] [0.0000 ; 0.0000] [-2.6617 ; 2.8778]
Mixtral-8x7B-Instruct-v0.1 [-3.4024 ; 2.0495] [-2.5194 ; 2.9893] [-2.4675 ; 3.0437] [-2.7556 ; 2.7884] [-2.8778 ; 2.6617] [0.0000 ; 0.0000]
Table 11: 95% Credible Intervals for posterior differences between model for attractor position - length
rephrase inspiration continue
rephrase [0.0000 ; 0.0000] [-4.1791 ; 1.3268] [-3.1691 ; 2.4195]
inspiration [-1.3268 ; 4.1791] [0.0000 ; 0.0000] [-1.7409 ; 3.8265]
continue [-2.4195 ; 3.1691] [-3.8265 ; 1.7409] [0.0000 ; 0.0000]
Table 12: 95% Credible Intervals for posterior differences between prompt for attractor position - length
B.4 Discontinuities and collapsing behavior
@@
Figure S7 shows the average length of text generated with the Mistral-7B model chain on the Continue task for five different seeds of the same story. We can observe several discontinuities in terms of the generated text length, i.e. at some iterations the length drastically increases or decreases. It is interesting to note that when the length decreases, it returns to the original value as before the first discontinuity. This suggests the existence of an attractor regarding this specific length. To better understand the cause of these discontinuities, Figure 13 shows examples of stories generated before and after those discontinuities (for the seed number three in figure S7). We can see that at generation 14 the model abruptly starts to generate many hashtags. It generates 283 hashtags, compared to 12 in the previous generation. At generation 45, we can see that the overall quality of the text decreased into generating solely hashtags and brief descriptions. This reduction in text quality is reminiscent of collapsing dynamics observed in iterative chains of LLMs, where each model was trained on the output of a previous one [Shumailov et al., 2023].